# RoTBench: A Multi-Level Benchmark for Evaluating the Robustness of Large Language Models in Tool Learning

### v2

@@ 4.2 Experimental Results
Figure 6: The means and standard deviations of our model’s performance in the five environments. “w/o” denotes that the stage is excluded from the RoTLLaMA training process.
@@
Further Studies
We compare the number of tool hallucinations for each LLM in all environments and find that our model has significantly fewer hallucinations compared to the GPT family of models (Table 6). This demonstrates the effectiveness of our method in mitigating interference from various sources of noise while accurately acquiring environmental information. It’s worth noting that the NexusRaven family of models, which relies on CodeLLaMA as a base, also exhibits low tool hallucinations, suggesting that utilizing code-based approaches for tool learning is a viable direction.
ToolLLaMA-2- NexusRaven- GPT- RoTLLaMA
7B-v1 7B-v2 13B-v1 13B-v2 3.5-turbo 4
53 65 6 0 50 23 3
5 Related Work
Analysis of Tool Learning
Given their extensive world knowledge and superior natural language understanding, researchers have made attempts to leverage LLMs for a wide range of everyday applications Ye et al. (2023). In order to push the boundaries of their capabilities, some scholars have proposed enhancing LLMs with external tools, which has gained widespread acceptance Schick et al. (2023); Tang et al. (2023). As research in this area has deepened, certain scholars have summarized the progress made in tool learning for LLMs Mialon et al. (2023); Qin et al. (2023a), sought to uncover developmental insights, and trained more specialized LLMs for tool learning based on these findings Qin et al. (2023b); Zhuang et al. (2023); Hao et al. (2023). Furthermore, recognizing the complexity of tool learning, some researchers have specialized in evaluating not only the outcomes of tool learning Huang et al. (2023) but also the entire process Chen et al. (2023d); Ye et al. (2024). However, it’s worth noting that all of these current efforts primarily consider LLMs’ tool usage in controlled environments, neglecting the inherent complexities of real-life scenarios. Therefore, we have undertaken an in-depth analysis of the robustness of LLMs in tool learning to advance research in a real-world context.
Robustness Testing of LLMs
Robustness is a critical factor in determining the stability of LLMs and plays a pivotal role in their practical deployment in real-life applications, which has garnered significant attention from scholars. In the early stages of research, some scholars conducted tests to assess the robustness of ChatGPT across various natural language processing tasks, highlighting the substantial room for improvement in the current robustness of LLMs Wang et al. (2023a); Chen et al. (2023c). Subsequently, other researchers specialized in creating benchmarks, such as PromptBench Zhu et al. (2023), to examine the consistency of LLM responses by introducing noise into the prompts. Given that tool learning is poised to extend the capabilities of LLMs and its outcomes can directly impact the state of the physical world Ye et al. (2024), it becomes imperative to thoroughly evaluate its robustness.
@@ 6 Conclusion
While we introduce a multi-level benchmark for evaluating the robustness of LLMs in tool learning and a training method aimed at increasing environmental diversity, our work does have some limitations. On one hand, our primary focus is on assessing the robustness of LLMs in a single tool-use round, and we do not delve into whether LLMs are able to self-correct their behavior in response to environmental feedback. However, we analyze the performance of GPT-4 based on the interaction trajectories in the first two rounds and find that this does not enhance model robustness. On the other hand, we focus our analysis on the noise of the tool names and parameters, and the information describing the default tools and parameters is accurate. This is because if there is an information inaccuracy in the tool description and parameter description, then the tool is completely unusable and not a noise that needs to be recognized.
Appendix A Examples for Noise Correction
Models Stage Query Noisy Part Model Output
GPT-3.5-turbo Tool Selection
I have a list of names: Maria, Juan, and Carlos. Can you predict their ages?
Tool: predOict_aTge
Description: Predicts the ages of one or more people given their names.
Parameters: …
Tool: predict_age
GPT-3.5-turbo Parameter Identification
I want to know what will be the output if we run these commands sequentially in bash: ‘cd /home/user/documents’, ‘ls -a.’
Tool: execute_bash_code
Description: …
Parameters: Nommands (Required)
Param Description: The command string to be executed.
Parameters: commands
GPT-4 Tool Selection
Is there any social event available which requires high accessibility and is free of cost?
Tool: get_activty_by_ye
Description: Find a random activity with a given type.
Parameters: …
Tool: get_activity_by_type
GPT-4 Parameter Identification
Get me quotes for symbols AAPL, MSFT, and GOOGL from US.
Tool: get_quotes
Description: …
Parameters: ymbols (Required)
Param Description: The value of symbol field returned in auto-complete endpoint. Separated by comma for multiple entities.
Parameters: symbols
Table 7: Examples for noise correction of GPT family of models.
Appendix B Prompt Template for Inference
System
First I will give you the task description, and your task start.
Then you will analyze your status now, then decide what to do next…
Desired format:
Thought: $\langle$ The $\rangle$
Action: $\langle$ The tool you decide to $\rangle$
Action Input: $\langle$ The parameters for the $\rangle$
Remember:
2. The action to take should be one of the given tools below.
{Tool Document}
Let’s Begin!
User
{Query}
Begin!
User
{Tool Document}
User Query: Question: {Query}
User
{Tool Document}
User Query: {Query}
Appendix C Prompt Template for Query Expansion
We use GPT-4 for query expansion based on prompt in Table 11.
System
Remember:
Here is the documentation of various tools: {Tool Document}
User
Please generate 12 diverse queries according to the documentation.
Examples:
{Examples}
Appendix D Prompt Template for Trajectory Generation
We use GPT-4 for trajectory generation based on prompt in Table 12.
System
Then you will analyze your status now, then decide what to do next…
Remember:
Let’s begin!
User
{Query}
Begin!
Appendix E Results in Different Scenarios
We show the performance of each model in different scenarios and document the results from Table 13 to Table 19.

### v3

@@ 6 Conclusion
While we introduce a multi-level benchmark for evaluating the robustness of LLMs in tool learning and a training method aimed at increasing environmental diversity, our work does have some limitations. On one hand, our primary focus is on assessing the robustness of LLMs in a single tool-use round, and we do not delve into whether LLMs are able to self-correct their behavior in response to environmental feedback. However, we analyze the performance of GPT-4 based on the interaction trajectories in the first two rounds and find that this does not enhance model robustness. On the other hand, While tool descriptions are undoubtedly crucial for understanding tools, our analysis centers on the noise present in tool names and parameters. This choice is driven by our discovery that LLMs’ comprehension of tools primarily relies on tool and parameter names rather than a nuanced understanding of the meanings conveyed in tool documentation. Within this framework, evaluating LLMs through RoTBench can effectively measure their tolerance to noise in these additional details, thus propelling research endeavors aimed at improving LLMs’ tool learning capabilities.
Acknowledgements
The authors wish to thank the anonymous reviewers for their helpful comments. This work was partially funded by National Natural Science Foundation of China (No. 62476061,62206057,62076069), Shanghai Rising-Star Program (23QA1400200), Natural Science Foundation of Shanghai (23ZR1403500), Program of Shanghai Academic Research Leader under grant 22XD1401100.
Appendix A Details of LLMs
A.1 Open-Source LLMs
ToolLLaMA-2-7B-v1
ToolLLaMA-2-7B-v1, developed by Tsinghua University, is a tool-oriented LLM that harnesses the power of 126,000 data samples, including more than 16,000 APIs, through supervised fine-tuning on LLaMA-2-7B-base. This enables ToolLLaMA-2-7B-v1 to effectively utilize various tools to meet diverse user requirements.
ToolLLaMA-2-7B-v2
ToolLLaMA-2-7B-v2 has undergone fine-tuning from LLaMA-2-7B-base, by assimilating an expansive dataset comprising over 120,000 solution paths and annotated chains of thought. To the best of our knowledge, this model stands as the most extensively trained tool-oriented LLM, utilizing the largest dataset and the broadest spectrum of tools among all available options.
NexusRaven-13B-v1
NexusRaven-13B-v1 is a tool-oriented model that underwent fine-tuning based on CodeLLaMA-13B. Distinguishing itself from prior models, NexusRaven-13B-v1 employs code nesting to invoke tools, generating the entire inference path simultaneously instead of following a step-by-step approach.
NexusRaven-13B-v2
NexusRaven-13B-v2 enhances the performance of NexusRaven-13B-v1 by generating single, nested, and parallel function calls in various complex scenarios. Additionally, NexusRaven-13B-v2 can generate inference paths for the function calls it creates, thereby improving overall generalization.
A.2 Closed-Source LLMs
GPT-3.5-turbo
GPT-3.5-turbo stands out as the most potent and cost-efficient model within the GPT-3.5 series. Tailored for conversations, it excels in comprehending and generating natural language. Furthermore, it exhibits strong tool invocation capabilities.
GPT-4
GPT-4 represents OpenAI’s most robust LLM, surpassing its predecessor in delivering safer and more beneficial responses. Additionally, GPT-4 offers formal support for multimodal inputs and has an expanded capability to address a broader spectrum of social requirements.
Appendix B Experimental Setup
Inference
In accordance with Ye et al. (2024a), we adopt the ReAct Yao et al. (2023) format for inference, employing a consistent prompt template for both the ToolLLaMA-2-7B family of models and the GPT family of models. However, as NexusRaven-13B fmaily of models utilize nested functions for output, we adhere to the guidelines outlined on their official website, which necessitate the use of a distinct set of template. [^8] The specific prompt can be found in Appendix F. Meanwhile, to evaluate human performance across environments with different noise levels, we enlist three university students. Each student receives identical tool documentation and task descriptions. Independently, they completes the questions and the average score derived from their responses served as the human performance benchmark.
Evaluation
We score the performance of LLMs and Human using the evaluation methods defined in Section 3.3. In this system, each data point is scored as 0 or 1 at each stage. This is because, in the context of tool learning, tool calls either succeed or fail, and even small errors can cause the entire process to fail. In particular, In the tool selection phase, an error in tool selection can lead to overall failure, independent of parameter accuracy. In the parameter identification phase, missing necessary parameters or wrong parameter selection can lead to failure. In the content filling phase, incorrect content input can lead to undesirable tool execution results.
Appendix C Results in Different Scenarios
We show the performance of each model in different scenarios and document the results from Table 6 to Table 12. From the results, we have the following observations.
The variance in average performance of LLMs across various study scenarios can be attributed to the relevance of specific features of available tools to each scenario.
For instance, in both application operations and personal life scenarios, LLMs may err due to the strict sequential order in which tools are called (e.g., obtaining parameter values for “list_properties” necessitates prior execution of “search_locations”).
It’s notable that the model’s perception of environmental complexity may diverge from human intentions.
For instance, in information retrieval scenarios, LLMs exhibit inferior average performance in the slight-level environment compared to the medium-level and heavy-level environments, primarily due to limitations in noise correction capabilities (Section 4.3).
Regarding the model itself, variations in training methods and data can lead to unexpected performances in certain scenarios.
For instance, ToolLLaMA-7B-v1 demonstrates a performance discrepancy between the clean-level and union-level environments in the application manipulation scenario, scoring 20 and 40, respectively. This disparity arises from its ability to perform better when only two tools are available alongside “ask_to_user” and “finish,” whereas GPT4 consistently prompts for API keys even when unnecessary.
@@ Appendix C Results in Different Scenarios
Appendix D Examples for Noise Correction
Models Stage Query Noisy Part Model Output
GPT-3.5-turbo Tool Selection I have a list of names: Maria, Juan, and Carlos. Can you predict their ages? Tool: predOict_aTge Description: Predicts the ages of one or more people given their names. Parameters: … Tool: predict_age
GPT-3.5-turbo Parameter Identification I want to know what will be the output if we run these commands sequentially in bash: ‘cd /home/user/documents’, ‘ls -a.’ Tool: execute_bash_code Description: … Parameters: Nommands (Required) Param Description: The command string to be executed. Parameters: commands
GPT-4 Tool Selection Is there any social event available which requires high accessibility and is free of cost? Tool: get_activty_by_ye Description: Find a random activity with a given type. Parameters: … Tool: get_activity_by_type
GPT-4 Parameter Identification Get me quotes for symbols AAPL, MSFT, and GOOGL from US. Tool: get_quotes Description: … Parameters: ymbols (Required) Param Description: The value of symbol field returned in auto-complete endpoint. Separated by comma for multiple entities. Parameters: symbols
Table 13: Examples for noise correction of GPT family of models.
Appendix E Further Studies about RoTTuning
We conduct additional comparative analysis to further validate the effectiveness of RoTTuning in improving the stability of LLMs in noisy environments.
Robust Generalization of RoTTuning
To validate the robust generalization of RoTTuning across different environments, we apply a single environment augmentation and compare the results to those without augmentation. As shown in Table  14, even when training RoTTuning with data from only one environment, it achieves superior performance in other environments, demonstrating its strong generalization capability.
Approaches w/o Augmentation w/ Aug.Slight w/ Aug.Medium w/ Aug.Heavy w/ Aug.Union
Tool Selection
Clean 74.29 70.48 72.38 75.24 71.43
Slight 65.24 71.90 62.38 69.05 64.29
Medium 61.90 68.57 65.71 70.95 66.67
Heavy 50.48 51.90 49.52 60.48 55.24
Union 40.00 53.33 51.43 53.33 55.24
Parameter Identification
Clean 60.95 57.14 59.05 59.05 60.95
Slight 47.14 53.81 46.19 48.10 46.19
Medium 42.86 51.90 48.57 48.57 52.38
Heavy 14.29 18.10 15.24 33.81 26.67
Union 21.90 32.38 28.57 31.43 36.19
Content Filling
Clean 45.71 43.81 48.57 44.76 42.86
Slight 31.90 40.00 31.90 35.24 30.95
Medium 30.48 38.10 36.67 36.67 38.57
Heavy 10.48 12.86 10.48 24.76 19.05
Union 12.38 19.05 17.14 21.90 27.62
Table 14: Performance of the LLMs trained by data augmented from single environment, compared with the model trained using LoRA without augmentation. The best performance in each environment is highlighted in bold.
The Number of Tool Hallucinations
We compare the number of tool hallucinations for each LLM in all environments and find that our model has significantly fewer hallucinations compared to the GPT family of models (Table 15). This demonstrates the effectiveness of our method in mitigating interference from various sources of noise while accurately acquiring environmental information. It’s worth noting that the NexusRaven family of models, which relies on CodeLLaMA Rozière et al. (2023) as a base, also exhibits low tool hallucinations, suggesting that utilizing code-based approaches for tool learning is a viable direction.
ToolLLaMA-2- NexusRaven- GPT- RoTLLaMA
7B-v1 7B-v2 13B-v1 13B-v2 3.5-turbo 4
53 65 6 0 50 23 3
Performance of RoTToolLLaMA
To confirm the robustness of our method for enhancing established tool-oriented LLMs, we proceed to fine-tune ToolLLaMA-2-7B using our generated trajectories and obtain RoTToolLLaMA. The corresponding results presented in Table 16 illustrate that our method’s fine-tuning significantly enhances the model’s tool learning capability across all stages, while also bolstering its overall robustness. For instance, across the three stages, our method demonstrates performance extremes of 12.33/13.33/9.53 in various environments, compared to ToolLLaMA-2-7B-v2’s 26.67/16.67/10.95. This further underscores the efficacy of our proposed approach.
Level Clean Slight Medium Heavy Union
$\mathbf{s_{TS}}$ 69.52 69.05 70.95 64.76 56.19
$\mathbf{s_{PI}}$ 52.38 45.24 50.95 40.95 39.05
$\mathbf{s_{CF}}$ 38.10 32.38 34.76 31.43 28.57
Table 16: The score in different stages (%) of RoTToolLLaMA in various Environments.
Appendix F Prompt Template for Inference
System
First I will give you the task description, and your task start.
Then you will analyze your status now, then decide what to do next…
Desired format:
Thought: $\langle$ The $\rangle$
Action: $\langle$ The tool you decide to $\rangle$
Action Input: $\langle$ The parameters for the $\rangle$
Remember:
2. The action to take should be one of the given tools below.
{Tool Document}
Let’s Begin!
User
{Query}
Begin!
User
{Tool Document}
User Query: Question: {Query}
User
{Tool Document}
User Query: {Query}
Appendix G Prompt Template for Query Expansion
We use GPT-4 for query expansion based on prompt in Table 20.
System
Remember:
Here is the documentation of various tools: {Tool Document}
User
Please generate 12 diverse queries according to the documentation.
Examples:
{Examples}
Appendix H Prompt Template for Trajectory Generation
We use GPT-4 for trajectory generation based on prompt in Table 21.
System
Then you will analyze your status now, then decide what to do next…
Remember:
Let’s begin!
User
{Query}
Begin!